# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v1

@@ 4.3 Greedy Client Center Selection & Domain Data Retrieval
Greedy client center selection. We will consider this problem from two aspects: 1) Select a client center that can represent the distribution of the local data. 2) To maximize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers. Firstly, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$. Then we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size, while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$. Specifically, for the $i$-th iteration, we filter $i$-top cluster centers that have the largest summed similarity with previously selected client centers and select the center of the rest largest cluster as the $i$-th client center. This procedure is repeated until have selected $N$ client centers.
Domain data retrieval. For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients. Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
In summary, FedDCA can be seen as a method for establishing a comprehensive training set that captures the essence of the domain and covers its distribution. Through the iterative selection of client centers that provide unique coverage, the algorithm ensures that the selected centers are representative of the domain and achieve great domain coverage. In addition, domain data retrieval exposes the model to a broader range of in-domain features, facilitating a more robust understanding of domain-specific tasks. This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, which ends up with performance degradation. By leveraging diverse examples across clients, the model can better generalize its learning, leading to enhanced performance on various tasks within the domain.
@@ 4.4 Heterogeneous Encoder with Feature Alignment
Due to the limited computational resources on the client side, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side. However, the output dimension of heterogeneous encoders may not be consistent. Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4, we use contrastive learning(Khosla et al. (2020)) and train $w_{p}$ on the public instructions. Specifically, $w_{p}$ is composed of two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$. For $\forall{x_{i}}\in D^{p}$, where $i\in\{1,2,\dots,\mathcal{N}\}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$. For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$. Let the batch size be $B$, then the training objective is defined as follows:
@@
Using a projector $w_{p}$ to achieve feature alignment is an effective strategy in FL for scenarios where clients and the server utilize encoders of different sizes or capacities. The projector maps the lower-dimensional feature representations $\mathcal{H}$ from the client-side encoder to the higher-dimensional space $\mathcal{H}^{\prime}$ of the server-side encoder. By presenting the projector with positive and negative pairs, $w_{p}$ learns to adjust the projections such that the feature representations of similar items are close to each other, while those of dissimilar items are distant. In conclusion, the heterogenous encoder setting allows the system to be more scalable and flexible based on different client capacities. As a result, the projector helps in creating a unified feature space where representations from diverse sources can be combined, and utilized collaboratively.
@@ 4.5 Discussions
Computation. Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 3 in the polynomial time and quite efficient. On the client side, the k-means algorithm(Vardakas & Likas (2023)) is $\mathcal{O}(|D_{k}^{l}|)$. On the server side, as mentioned in Section 4.2, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$. Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$. For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset, thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the hetergenous encoder setting FedDCA further reduces the computation overhead on the client side.
Communication. In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server. Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, which size is $N_{k}^{p}$. In addition, in the fine-tuning stage, FedDCA follows the standard FedIT procedure, which exchanges the LoRA parameters $\Delta\phi_{k}$ between clients and the server. Specifically, for Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million. Compared with the number of the frozen pretrained LLM parameters, the communication for LoRA tuning is quite efficient.
Privacy. The proposed method FedDCA does not violate the client’s privacy. Compared with other FedIT methods(Zhang et al. (2024d); Ye et al. (2024b)), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption(Acar et al. (2018)), which allows the server directly computates on ciphertext for matrix multiplication for dense retrieval.
@@ 5 Experiments
To show the effectiveness of FedDCA, we conduct extensive experiments on various domains and compare FedDCA with several baselines. This section is organized as follows: First we give details of experimental setup in Section 5.1, then we report the performance of FedDCA and compare it with several baselines in Section 5.2, and next we evaluate the computational cost between the FedDCA and FedDCA in Section 5.3. Finally, we do privacy evaluations of FedDCA in Section 5.4. For more details and results please see in Appendix A.
@@ 5.1 Experimental Setup
Dataset and Evaluation Metrics. To evaluate the performance of FedDCA, we conduct experiments on the four domains: code, medical, financial and mathematcial. The detail of constructing the public dataset is described in Appendix A.5. We choose HumanEval (code, H-Eval for short)(Chen et al. (2021)), MMLU-Med (medical, abbreviated as M-Med)(Hendrycks et al. (2021)), FPB (financial), FiQA (financial), TFNS (financial)(Yang et al. (2023a)), GSM8K (mathematical)(Cobbe et al. (2021)) for evaluation. HumanEval is evaluated using Pass@1. For MMLU-Med, FPB, FiQA and TFNS, we use accuracy as evaluation metric and exact match (EM) is used for GSM8K. Please see Appendix A.5 for more dataset details.
Baselines. We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented. Zero-shot inference and FedIT, which composes of four widely used FL methods, including FedAvg(McMahan et al. (2017)), FedProx(Li et al. (2020)), SCAFFOLD(pmlr-v119-karimireddy20a) and FedAvgM(Hsu et al. (2019)). 2) Augmented. Random sampling, direct retrieval, LESS(Xia et al. (2024)), Self-Instruct(Wang et al. (2023c)). Additionally, we report FedDCA’s performance with different FL strategies in Table 2. Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound. Direct retrieval is described in Appendix A.3. Self-Instruct augments instruction from a generative way through GPT-3.5-turbo(Sun et al. (2023)). Base on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
@@ 5.2 Performance Analysis
Performance. We evaluate the performance of FedDCA and compare it with several baselines on four domains (see training details in Appendix A.6). As shown in Table 2, FedDCA outperforms the other nine baselines in all domains. In particular, FedDCA+FedProx achieves the best performance in most domains, with a substantial improvement from at least 5.61% to the maximum of 14.63% over other baselines on the average performance. In addition, to reduce the computation overhead of clients, FedDCA attempts to utilize heterogeneous encoders. We see that although FedDCA has a performance drop of 1.67% than FedDCA, it still outperforms other baselines in the average performance. Furthermore, we report FedDCA’s performance with different FL strategies, we can see that no FL methods can keep the leading position in all domains. In specific, FedProx and SCAFFOLD FL strategies perform better in the average performance. Overall, the result shows the effectiveness of FedDCA and FedDCA.
Table 2: Performance(%) of FedDCA, FedDCA and other nine baselines on various domains. We report FedDCA with different FL strategies in the last four rows and FedDCA+FedProx achieves the best average performance across four domains.
@@
Furthermore, Table 2 can be divided into two parts: unaugmented methods and augmented methods. We can observe that FedIT methods perform well in the code domain, even better than most augmented methods. This could be attributed to two reasons: 1) The code domain is more about following a certain paradigm. 2) As the local data is few, so compared with the same epoch and batch size, the unaugmented methods learn the same data more times, which is kind of not a fair setting. However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, which further demonstrates the effectiveness of FedDCA.
Efficiency. We see that direct retrieval is the best baseline in the average performance. However, it does not consider the cross-client domain coverage, resulting in an overlapping retrieval data distribution. Self-Instruct(Wang et al. (2023c)) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT(Zhang et al. (2024d)) attempts to augment local data by leveraging the pre-trained Llama2 model(Li & Flanigan (2023)) to perform self-instruct manner instruction generation during training. However, there exist two concerns: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly in terms of both time and computational resources. Overall, generating instructions through a self-instruct manner is not a satisfactory method for the FedDIT scenario. LESS(Xia et al. (2024)) represents the augmentation methods through gradient feature retrieval, however, the performance of LESS is not satisfactory. What’s worse, it assumes success to the validation set which is not always available and requires a warmup process on the public dataset first and then computes the gradient of both the public data and the validation set, which is computationally expensive when the size of the public dataset is enormous. In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows the efficiency of FedDCA.
@@
Domain coverage. The domain coverage of each method is shown in Table 3. Also, we show the visualization of the domain coverage across different instruction augmentation methods in Appendix A.10. For the augmented methods, it can be observed that the higher domain coverage in a particular domain correlates with the better performance of the method in that domain. Specifically, given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all four domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement. In addition, as FedDCA utilizes a smaller encoder on the client side for computational efficiency, it sacrifices some semantic precision, resulting in a 2.89% averaged domain coverage drop in relative. However, FedDCA still surpasses the best baseline by 1.79% on average. In brief, the result proves the effectiveness of FedDCA and FedDCA in domain coverage augmentation.
@@ 5.3 Computation Analysis
Evaluation setup. The encoders used for FedDCA and FedDCA on the client side are bge-large-en-v1.5 and all-MiniLM-L6-v2 respectively (detailed in Appendix A.6). Then we show the model size and the time overhead of encoding the public instructions in Table 4
We can observe that in comparison with FedDCA, there is a large computational and time overhead advantage with FedDCA, while achieves an acceptable performance (shown in Table 2), which is accomplished by implementing a learnable projector on the server side. This projector maps the lower-dimensional client-side feature to the higher-dimensional space of the server-side encoder output. Consequently, the projector facilitates the creation of a unified feature space where representations from diverse sources can be integrated and used collaboratively.
@@ 5.4 Privacy Analysis
Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks(Carlini et al. (2021); Zhang et al. (2024c)), which utilizes the autoregression nature of LLM to extract information from the memory(Xu et al. (2024)).
@@
Evaluation setup. We focus on a single client’s fine-tuning process, which has 100 local data and adopts FedDCA to do the instruction augmentation with different amounts of public data from 1000 to 5000. We set up 10 clients with full participation for 10 rounds. Specifically, we record every round’s ROUGE-L score(Lin (2004)) of the client $c_{0}$. We use the designed prompt to extract the instruction that LLM memorizes, which is shown in Appdix A.8. Specifically, for each setting, we repeat the memory extraction 100 times and report the average ROUGE-L score computed based on each generated instruction and all the local data’s instructions to evaluate the privacy-preserving capability of different amounts of public data.
Figure 3: ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
Results. We report the average ROUGE-L score of base-data-only fine-tuning and various amounts of public data fine-tuning in Table 5. The result shows that at the same training round, there is no significant correlation between the public data ratio and the privacy-preserving capability. In addition, after 10 rounds of FedDIT, the base-data-only fine-tuning’s average ROUGE-L score all exceeds the instruction-augmented settings in four domains. Specifically, base-data-only fine-tuning exceeds augmented settings by 6.69%, 5.55%, 5.55%, and 6.16% on average, relatively.
The trend of the ROUGE-L score of base-data-only and augmented fine-tuning of each domain per round is shown in Figure 3. In specific, the augmented fine-tuning uses 5000 public data. We can see that the ROUGE-L score of augmented settings in each domain firstly increases and then starts decreasing or converging to a certain range. The base-data-only setting keeps the rising trend, especially the code domain, which indicates that as the number of training rounds increases, the base-data-only fine-tuning will remember more privacy information, while the risk of privacy leakage of the augmented fine-tuning will gradually decrease or converge.

### v3

@@ 4.2 Greedy Client Center Selection & Domain Data Retrieval
Greedy client center selection. We will consider this problem from two aspects: 1) Select a client center that can represent the distribution of the local data. 2) To optimize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers. First, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$. Then, we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$. Specifically, for the $i$-th iteration, we filter $i$-top cluster centers that have the largest summed similarity with previously selected client centers and select the center of the rest largest cluster as the $i$-th client center. This procedure is repeated until we have selected $N$ client centers.
Domain data retrieval. For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients. Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
In summary, FedDCA establishes a comprehensive training set that captures the essence and distribution of the domain by iteratively selecting client centers with unique coverage increments. This ensures that selected centers are representative and achieve broad domain coverage. Additionally, domain data retrieval exposes the model to a wider range of in-domain features, enhancing its understanding of domain-specific tasks. This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, ultimately improving performance across various tasks within the domain.
@@ 4.3 Heterogeneous Encoder with Feature Alignment
For the client-side computational efficiency and system scalability, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side. However, the output dimension of heterogeneous encoders may not be consistent. Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4. We use contrastive learning (Khosla et al., 2020) and train $w_{p}$ on the public instructions. Specifically, $w_{p}$ comprises two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$. For $\forall{x_{i}}\in D^{p}$, where $i\in\{1,2,\dots,\mathcal{N}\}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$. For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$. Let the batch size be $B$, then the training objective is defined as follows:
@@
The heterogeneous encoder setting, by employing a projector $w_{p}$ for feature alignment, proves effective in FL contexts where clients and the server use different encoder sizes or capacities. This approach facilitates mapping lower-dimensional client features $\mathcal{H}$ to the higher-dimensional server space $\mathcal{H}^{\prime}$ through a strategic training of $w_{p}$ with both positive and negative pairs. It adeptly aligns similar feature representations closely while positioning dissimilar ones further apart. As a result, the heterogeneous encoder setting not only enhances system scalability and flexibility but also facilitates the creation of a unified feature space that effectively integrates and utilizes diverse representations collaboratively.
@@ 4.4 Discussions
Computation. Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 3 in the polynomial time and is quite efficient. On the client side, the k-means algorithm (Vardakas & Likas, 2023) is $\mathcal{O}(|D_{k}^{l}|)$. On the server side, as mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$. Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$. For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset. Thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
Communication. In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server. Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, whose size is $N_{k}^{p}$. In addition, in the fine-tuning stage, FedDCA follows the standard FedIT procedure, which exchanges the LoRA parameters $\Delta\phi_{k}$ between clients and the server. Specifically, for the Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million. Compared with the number of frozen pre-trained LLM parameters, the communication for LoRA tuning is quite efficient.
Privacy. The proposed method, FedDCA, does not violate the client’s privacy. Compared with other FedIT methods (Zhang et al., 2024d; Ye et al., 2024b), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption (Acar et al., 2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
@@ 5 Experiments
To demonstrate the effectiveness of FedDCA and its variant FedDCA, we conduct extensive experiments across various domains and with several baselines. For additional details and results, please refer to Appendix A.
@@ 5.1 Experimental Setup
Dataset and Evaluation Metrics. To evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical. The detail of constructing the public dataset is described in Appendix A.5. For evaluation, we select a range of datasets, including HumanEval (H-Eval for short) for coding (Chen et al., 2021), MMLU-Med (abbreviated as M-Med) for medical (Hendrycks et al., 2021), and GSM8K for mathematics (Cobbe et al., 2021). Additionally, financial datasets such as FPB, FiQA, and TFNS (Yang et al., 2023a) are utilized. HumanEval is evaluated using Pass@1. For MMLU-Med, FPB, FiQA, and TFNS, we use accuracy as the evaluation metric. While GSM8K is evaluated using exact match (EM). Please see Appendix A.5 for more dataset details.
Baselines. We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a) and FedAvgM (Hsu et al., 2019). 2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2023c). Additionally, we report FedDCA’s performance with different FL strategies in Table 2. Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound. Direct retrieval is described in Appendix A.3. Self-Instruct augments instruction in a generative way through GPT-3.5-turbo (Sun et al., 2023). Based on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
@@ 5.2 Performance Analysis
Performance. We evaluate the performance of FedDCA and compare it with several baselines on four domains (see training details in Appendix A.6). As shown in Table 2, FedDCA outperforms the other nine baselines in all domains. In particular, FedDCA+FedProx achieves the best performance in most domains, with a substantial improvement from at least 5.61% to the maximum of 14.63% over other baselines on the average performance. In addition, to reduce the computation overhead of clients, FedDCA attempts to utilize heterogeneous encoders. We see that although FedDCA has a performance drop of 1.85% than FedDCA, it still outperforms other baselines in the average performance. Furthermore, we report FedDCA’s performance with different FL strategies, and we can see that no FL method can keep the leading position in all domains. In specific, FedProx and SCAFFOLD FL strategies perform better in the average performance. Overall, the result shows the effectiveness of FedDCA and FedDCA.
Table 2: Performance(%) of FedDCA, FedDCA and other nine baselines on various domains. We report FedDCA with different FL strategies in the last four rows, and FedDCA+FedProx achieves the best average performance across four domains.
@@
Additionally, Table 2 can be divided into two parts: unaugmented methods and augmented methods. We can observe that FedIT methods perform well in the code domain, even better than most augmented methods. This could be attributed to two reasons: 1) The code domain is more about following a certain paradigm. 2) As the local data is few, so compared with the same epoch and batch size, the unaugmented methods learn the same data more times, which is kind of not a fair setting. However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, further demonstrating the effectiveness of FedDCA.
Efficiency. We see that direct retrieval is the best baseline in the average performance. However, it does not consider cross-client domain coverage, resulting in an overlapping retrieved data distribution. Self-Instruct (Wang et al., 2023c) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT (Zhang et al., 2024d) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training. However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly regarding both time and computational resources. Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario. LESS (Xia et al., 2024) represents the augmentation methods through gradient feature retrieval. However, its performance is underwhelming. Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data. This process becomes computationally burdensome with large public datasets. In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows its efficiency.
@@
Domain coverage. Each method’s domain coverage is shown in Table 3. Additionally, visualization of domain coverage across different strategies can be found in Appendix A.10. For augmented methods, a correlation is evident between higher domain coverage in a specific domain and improved performance of the method within that domain. Given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement. Furthermore, while FedDCA employs a smaller encoder on the client side to enhance computational efficiency, this leads to a slight compromise in semantic precision, observed as a relative drop of 2.89% in average domain coverage. Nevertheless, FedDCA still outperforms the best baseline by an average of 2.78%. In brief, the result proves the effectiveness of FedDCA and FedDCA in domain coverage augmentation.
@@ 5.3 Computation Analysis
Evaluation setup. The encoders used for FedDCA and FedDCA on the client side are bge-large-en-v1.5 and all-MiniLM-L6-v2 respectively (detailed in Appendix A.6). Then, we show the model size and the time overhead of encoding the public instructions in Table 4
Compared to FedDCA, FedDCA has significant computational and time advantages while maintaining acceptable performance (shown in Table 2). This is accomplished by employing heterogeneous encoders and configuring a projector on the server to achieve dimensional mapping.
@@ 5.4 Privacy Analysis
Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024c), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
@@
Evaluation setup. We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1000 to 5000 public instructions. We set up 10 clients with full participation for 10 rounds. Specifically, we record the average ROUGE-L score (Lin, 2004) of client $c_{0}$ for each round. The designed prompt to extract instructions memorized by the LLM is detailed in Appendix A.8. For each setting, we repeat memory extraction 100 times and report the average ROUGE-L score based on each generated instruction and all local data instructions to evaluate the privacy-preserving capability of different public data amounts. Specifically, denoting the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$. The calculation is defined as $\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
Figure 3: Average ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
Results. We report the average ROUGE-L scores for base-data-only and various public data fine-tuning in Table 5. The results show no significant correlation between the public data ratio and privacy-preserving capability in the same training round. After 10 rounds of FedDIT, base-data-only fine-tuning outperforms the instruction-augmented settings in all four domains by 6.69%, 5.55%, 5.55%, and 6.16% on average.
Figure 3 shows the average ROUGE-L score trends per round for each domain, where augmented fine-tuning uses 5000 public data. Initially, the ROUGE-L scores for augmented settings increase, then decrease or converge, while the base-data-only scores continue to rise, especially in the code domain. This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.